OpenStudy (anonymous):

What is the x-intercept for this linear equation? 1/2x-2/3y=-6

jimthompson5910 (jim_thompson5910):

Plug in y = 0 and solve for x

jimthompson5910 (jim_thompson5910):

1/2x-2/3y=-6 1/2x-(2/3)(0)=-6 1/2x = -6 Can you take it from here?

OpenStudy (anonymous):

I got it!

jimthompson5910 (jim_thompson5910):

that's great :)

OpenStudy (anonymous):

Yes! Thanks, you always help me a lot.

jimthompson5910 (jim_thompson5910):

I'm glad I'm helping you out.